# Analytic Linearization and Holomorphic Extensions of Proper Groupoids

### v2

@@ 1. Introduction
Proper Lie groupoids are known to be linearizable around their orbits and, more generally, around saturated submanifolds. This was conjectured by Alan Weinstein [20, 21] and later proved in full generality in [6, 24] (see also [18, 7]). All these works were carried out in the smooth category. In this paper, we show that analogous linearization results hold in the analytic category. We apply these results to study holomorphic extensions of analytic algebroids and groupoids.
@@
It is not difficult to see that if an analytic groupoid admits an analytic 2-metric, then the exponential map yields an analytic linearization around an orbit. In this paper, we show that an analytic s-proper Lie groupoid admits such an analytic 2-metric. The construction of a 2-metric relies on an averaging procedure adapted to the analytic setting, which in turn requires the existence of an analytic Haar density. In Appendix A, we establish that any s-proper analytic groupoid admits such a density. Hence, the exponential map provides an analytic linearization, yielding our first main result:
@@
Every s-proper analytic groupoid admits an analytic linearization around any invariant submanifold.
Let us turn now to holomorphic extensions. By a holomorphic extension of an analytic algebroid $A\Rightarrow M$ we mean a holomorphic Lie algebroid $A_{\mathbb{C}}\Rightarrow M_{\mathbb{C}}$ together with a totally real algebroid embedding $\iota:A\to A_{\mathbb{C}}$ (cf. Definition 4.3). It is not hard to see that any analytic Lie algebroid admits a holomorphic extension, that the germ of this extension is unique, and that it satisfies a universal property. We discuss these and other results about holomorphic extensions of analytic algebroids in Section 4.
At the groupoid level, the notion of holomorphic extension is more subtle. In Section 5, we introduce a notion holomorphic extension of an analytic groupoid which extends both the notion of complexification of a manifold and of holomorphic extension of a Lie group. Note that an analytic groupoid may not have a holomorphic extension. This is already the case for Lie groups. However, by a classical result of Chevalley, every compact Lie group admits a holomorphic extension, and our second main theorem extends this fact.
@@
Every analytic s-proper groupoid admits a holomorphic extension.
@@
  • Given an integrable analytic algebroid, does it admit a holomorphic extension that is also integrable?
@@
If an analytic algebroid $A$ admits an s-proper integration then $A$ admits a holomorphic extension which is integrable.
@@
Acknowledgments. We would like to thank Ana Balibanu, Matias del Hoyo, Brent Pym, Florian Zeiser, Maarten Mol for several valuable discussions that helped shape the ideas in this paper. We are in debt to David Martinez Torres for many comments and suggestions on an earlier draft of this paper.
@@ 2. Analytic 2-metrics
Convention. Through out this paper we assume that Lie groupoids are Hausdorff and source connected.
@@ 2.1. Analytic Riemannian submersions
$\mathrm{d}_{e}p:(\operatorname{ker}\mathrm{d}p)_{e}^{\perp}\xrightarrow{\raisebox{-1.99997pt}[0.0pt][0.0pt]{\smash{$\sim$}}}T_{b}B$
@@ 2.2. Transversely invariant analytic metrics
    proper if the map $(s,t):\mathcal{G}\times\mathcal{G}\rightarrow M$ is a proper map,
@@ 2.4. Averaging
It is clear that $\tilde{\eta}$ is analytic so it suffices to show that for any orbit $\mathcal{O}\subset E$, the conormal representation $\mathcal{N}^{*}_{\mathcal{O}}(\theta):\mathcal{G}\curvearrowright(T\mathcal{O})^{\circ}$ acts by isometries with respect to the metric $\tilde{\eta}^{*}|_{(T\mathcal{O})^{\circ}\times T\mathcal{O})^{\circ}}$. According to Theorem 2.9, for any $\alpha\in(T\mathcal{O})^{\circ}$, one has
@@ 2.5. Existence of analytic 2-metrics
Since the graph is commutative, we have
@@ 3. Analytic linearization of s-proper groupoids
Let $\mathcal{G}\rightrightarrows M$ be an analytic Lie groupoid and let $S\subset M$ be an invariant submanifold. Then $\mathcal{G}$ can be analytically invariantly linearizable around $S$.
@@ 4.1. Totally real submanifolds
    For any $x\in M$, there is a chart $U$ for $M_{\mathbb{C}}$ centered at $x$, such that
    $M\cap U=\{(x_{1},\dots,x_{n})\in\mathbb{C}^{n}\mid x_{i}\in\mathbb{R}\quad i=1,\dots,n\}$
@@
Notice that the germ of the holomorphic extension $\phi^{*}$, in the last part of the statement, is unique: if $\phi^{*}_{k}:V_{k}\to Y$, $k=1,2$, are two holomorphic extensions of $\phi:M\to Y$, then there is exists an open $V\subset V_{1}\cap V_{2}$ containing $M$ such that $\phi^{*}_{1}|_{V}=\phi^{*}_{2}|_{V}$.
@@ 4.2. Definition of holomorphic extension
Under the assumptions of this definition, one can also calls $A\Rightarrow M$ a real form of $A_{\mathbb{C}}\Rightarrow M_{\mathbb{C}}$. However, at the groupoid level these two notions become distinct. For this reason, we will not use this term.
@@ 6. Holomorphic extensions of s-proper groupoids
It is shown in [15] that every proper Lie groupoid admits a compatible analytic structure.
@@
    It is enough to show that a neighborhood of of the identity section $M\subset U\subset\mathcal{G}$ has a compatible analytic structure making $U$ a local analytic groupoid. Indeed, such a neighborhood generates $\mathcal{G}$, so one obtains a compatible analytic structure in the whole of $\mathcal{G}$.
@@ 6.3. Gluing the local holomorphic extensions
$\psi^{*}_{ij}:\mathcal{G}_{i}^{*}|_{T^{*}_{ij}}\xrightarrow{\raisebox{-1.99997pt}[0.0pt][0.0pt]{\smash{$\sim$}}}\mathcal{G}_{j}^{*}|_{T^{*}_{ij}},\quad\psi^{*}_{ji}=(\psi^{*}_{ij})^{-1}.$
@@
$\psi^{*}_{ij}:\mathcal{G}_{i}^{*}|_{W_{i}\cap W_{j}}\xrightarrow{\raisebox{-1.99997pt}[0.0pt][0.0pt]{\smash{$\sim$}}}\mathcal{G}_{j}^{*}|_{W_{i}\cap W_{j}},\quad\psi^{*}_{ji}=(\psi^{*}_{ij})^{-1}.$

### v3

@@ 1. Introduction
Proper Lie groupoids are known to be linearizable around their orbits and, more generally, around saturated submanifolds. This was conjectured by Alan Weinstein [22, 23] and later proved in full generality in [7, 26] (see also [20, 8]). All these works were carried out in the smooth category. In this paper, we show that analogous linearization results hold in the real analytic category. We apply these results to study holomorphic extensions of real analytic algebroids and groupoids.
@@
It is not difficult to see that if a real analytic groupoid admits an analytic 2-metric, then the exponential map yields an analytic linearization around an orbit. In this paper, we show that a real analytic s-proper Lie groupoid admits such an analytic 2-metric. The construction of a 2-metric relies on an averaging procedure adapted to the analytic setting, which in turn requires the existence of an analytic Haar density. In Appendix A, we establish that any s-proper real analytic groupoid admits such a density. Hence, the exponential map provides an analytic linearization, yielding our first main result:
@@
Every s-proper real analytic groupoid admits an analytic linearization around any invariant submanifold.
Note that it is shown in [17] that every proper Lie groupoid admits a compatible real analytic structure.
Theorem 1 can be applied to promote known linearization results from the smooth to the real analytic category. One example is the following real analytic version of [10, Theorem 8.6]:
Corollary 1.
Let $(M,\pi)$ be a real analytic Poisson manifold and $S\subset M$ a Poisson submanifold. If $T^{*}_{S}M$ is integrable by a compact, Hausdorff, Lie groupoid whose source fibers have trivial 2nd de Rham cohomology, then $(M,\pi)$ is analytically linearizable around $S$.
Let us turn now to holomorphic extensions. By a holomorphic extension of a real analytic algebroid $A\Rightarrow M$ we mean a holomorphic Lie algebroid $A_{\mathbb{C}}\Rightarrow M_{\mathbb{C}}$ together with a totally real algebroid embedding $\iota:A\to A_{\mathbb{C}}$ (cf. Definition 4.3). It is not hard to see that any real analytic Lie algebroid admits a holomorphic extension, that the germ of this extension is unique, and that it satisfies a universal property. We discuss these and other results about holomorphic extensions of real analytic algebroids in Section 4.
At the groupoid level, the notion of holomorphic extension is more subtle. In Section 5, we introduce a notion holomorphic extension of a real analytic groupoid which extends both the notion of complexification of a manifold and of holomorphic extension of a Lie group. Note that a real analytic groupoid may not have a holomorphic extension. This is already the case for Lie groups. However, by a classical result of Chevalley, every compact Lie group admits a holomorphic extension, and our second main theorem extends this fact.
@@
Every real analytic s-proper groupoid admits a holomorphic extension.
@@
  • Given an integrable real analytic algebroid, does it admit a holomorphic extension that is also integrable?
@@
If a real analytic algebroid $A$ admits an s-proper integration then $A$ admits a holomorphic extension which is integrable.
@@
Every real analytic Poisson manifold of s-proper type admits an extension to an integrable holomorphic Poisson manifold.
@@
Acknowledgments. We would like to thank Ana Balibanu, Matias del Hoyo, Brent Pym, Florian Zeiser, and Maarten Mol for several valuable discussions that helped shape the ideas in this paper. We are in debt to David Martinez Torres for many comments and suggestions on an earlier draft of this paper.
@@ 2. Analytic 2-metrics
Convention. Throughout this paper we assume that Lie groupoids are Hausdorff and source connected. Also, by an analytic algebroid/groupoid we means a real analytic one.
@@ 2.1. Analytic Riemannian submersions
$\mathrm{d}_{e}p:(\operatorname{ker}\mathrm{d}p)_{e}^{\perp}\xrightarrow{\raisebox{-1.59442pt}[0.0pt][0.0pt]{\smash{$\sim$}}}T_{b}B$
@@ 2.2. Transversely invariant analytic metrics
    proper if the map $(s,t):\mathcal{G}\rightarrow M\times M$ is proper,
@@ 2.4. Averaging
According to Lemma A.5, $\tilde{\eta}$ is analytic, so it suffices to show that for any orbit $\mathcal{O}\subset E$, the conormal representation $\mathcal{N}^{*}_{\mathcal{O}}(\theta):\mathcal{G}\curvearrowright(T\mathcal{O})^{\circ}$ acts by isometries with respect to the metric $\tilde{\eta}^{*}|_{(T\mathcal{O})^{\circ}\times T\mathcal{O})^{\circ}}$. According to Theorem 2.9, for any $\alpha\in(T\mathcal{O})^{\circ}$, one has
@@ 2.5. Existence of analytic 2-metrics
Since the diagram is commutative, we have
@@ 3. Analytic linearization of s-proper groupoids
Let $\mathcal{G}\rightrightarrows M$ be an analytic s-proper Lie groupoid and let $S\subset M$ be an invariant submanifold. Then $\mathcal{G}$ can be analytically invariantly linearizable around $S$.
This result can be applied to obtain analytic versions of known linearization results in the smooth category. For example, in Poisson geometry, one obtains the following analytic version of [10, Theorem 8.6]:
Corollary 3.4.
Let $(M,\pi)$ be an analytic Poisson manifold and $S\subset M$ a Poisson submanifold. If $T^{*}_{S}M$ is integrable by a compact, Hausdorff, Lie groupoid whose source fibers have trivial 2nd de Rham cohomology, then $(M,\pi)$ is analytically linearizable around $S$.
Proof.
Theorem 8.2 in [10] states that a Poisson manifold $(M,\pi)$ is linearizable around a Poisson submanifold $S$ if and only if the algebroid $T^{*}M$ is linearizable around $S$. The proof in [10] holds in the analytic category, so all we need to show is that the algebroid $T^{*}M$ is analytically linearizable around $S$. If we show that $(M,\pi)$ is integrable by an analytic s-proper Lie groupoid on an invariant neighborhood $U$ of $S$, then Theorem 3.3 shows that it is analytically linearizable around $S$, so its algebroid will be analytically linearizable around $S$, and the proof will be completed.
Corollary 8.9 in [10] shows that the assumptions made in the statement imply that there is an invariant neighborhood $U$ of $S$ such that $(U,\pi|_{U})$ is integrable by an s-proper Lie groupoid. But this groupoid has a unique compatible real analytic structure, since it integrates an analytic algebroid: this follows either by observing that the exponential map gives a real analytic structure in a neighborhood of the identity section and any such neighborhood generates the whole groupoid (Proposition 13.7 in [6]) or, alternatively, by using the result of [17] which shows that every proper groupod has a compatible analytic structure. ∎
@@ 4.1. Totally real submanifolds
    For any $x\in M$, there is a chart $U$ for $M_{\mathbb{C}}$ centered at $x$, such that MU={(x_1,…,x_n)C^n x_iR i=1,…,n}
@@
Notice that the germ of the holomorphic extension $\phi^{*}$, in the last part of the statement, is unique: if $\phi^{*}_{k}:V_{k}\to Y$, $k=1,2$, are two holomorphic extensions of $\phi:M\to Y$, then there exists an open $V\subset V_{1}\cap V_{2}$ containing $M$ such that $\phi^{*}_{1}|_{V}=\phi^{*}_{2}|_{V}$.
@@ 4.2. Definition of holomorphic extension
Under the assumptions of this definition, one can also call $A\Rightarrow M$ a real form of $A_{\mathbb{C}}\Rightarrow M_{\mathbb{C}}$. However, at the groupoid level these two notions become distinct. For this reason, we will not use this term.
@@ 5.2.4. Bundle of compact Lie groups
5.2.5. Holomorphic extension of compact Lie group actions
Let $G$ be a compact Lie group acting analytically on an analytic manifold $X$. Let $G_{\mathbb{C}}$ denote the complexification of $G$. By [12], the action extends to a holomorphic action of $G_{\mathbb{C}}$ on the complexification $X_{\mathbb{C}}$ of $X$. Consequently, the holomorphic action groupoid $G_{\mathbb{C}}\ltimes X_{\mathbb{C}}$ provides a holomorphic extension of the action groupoid $G\ltimes X$.
@@ 6. Holomorphic extensions of s-proper groupoids
    It is enough to show that a neighborhood of the identity section $M\subset U\subset\mathcal{G}$ has a compatible analytic structure making $U$ a local analytic groupoid. Indeed, such a neighborhood generates $\mathcal{G}$, so one obtains a compatible analytic structure in the whole of $\mathcal{G}$.
@@ 6.3. Gluing the local holomorphic extensions
$\psi^{*}_{ij}:\mathcal{G}_{i}^{*}|_{T^{*}_{ij}}\xrightarrow{\raisebox{-1.59442pt}[0.0pt][0.0pt]{\smash{$\sim$}}}\mathcal{G}_{j}^{*}|_{T^{*}_{ij}},\quad\psi^{*}_{ji}=(\psi^{*}_{ij})^{-1}.$
@@
$\psi^{*}_{ij}:\mathcal{G}_{i}^{*}|_{W_{i}\cap W_{j}}\xrightarrow{\raisebox{-1.59442pt}[0.0pt][0.0pt]{\smash{$\sim$}}}\mathcal{G}_{j}^{*}|_{W_{i}\cap W_{j}},\quad\psi^{*}_{ji}=(\psi^{*}_{ij})^{-1}.$